The histogram below gives the length of service of members of the Department of English at a particular university. The classes, in years of service, are 0−−4.9, 5−−9.9, etc., and the vertical axis represents the number of faculty.

(a) What percent of the department faculty have 30 or more years of service?
Answer:  %

(b) If a member of the department is chosen at random to serve on a university committee, what is the probability (in decimal form) that the chosen representitive will have between 5 and 35 years of service?
Answer:

(c) What is the probability the representative above will have less than 10 years of service given that the person has less than 30 years of service?
Answer

English Faculty 9 8 7 6 5 T I 4 3 2 1 0 5 10 15 20 25 30 35 40 45 Years of Service
Answer #1

(a) Percent = (4 + 9 + 4)/(7 + 5 + 6 + 1 + 5 + 9 + 4 + 9 + 4)*100 = 34%

(b) Probability = (5 + 6 + 1 + 5 + 9 + 4)/(7 + 5 + 6 + 1 + 5 + 9 + 4 + 9 + 4) = 0.6

(c) Probability = (7 + 5)/(7 + 5 + 6 + 1 + 5 + 9) = 0.3636

Please give me a thumbs-up if this helps you out. Thank you!